OpenStudy (anonymous):

Find the area of a rectangular window measuring 3/4 meter by 5/8 meter

OpenStudy (campbell_st):

the area is \[\frac{3 \times 5}{4 \times 8}\] then answer will have units of m^2

OpenStudy (anonymous):

m^2? what are you mean? mine asnwer is 15/32 Am i right or wrong?